Your employer has promised to give you a $5,000 bonus after you have been working for him for 5 years. What is the present value of this bonus if the proper discount rate is 8%?
Answer #1

P = F / (1+i)^t

F = 5000, i = 8%, t = 5 yrs

P = 5000 / (1+0.08)^5

= 5000 / 1.08^5

= 5000 / 1.469328

= 3402.91